hotcheetos:

The number line shows the distance in meters of two jellyfish, A and B, from a predator located at point X: A horizontal number line extends from negative 3 to positive 3. The point A is at negative 1.5, the point 0 is labeled as X, and the point labeled B is at 0.5. Write an expression using subtraction to find the distance between the two jellyfish. (5 points) Show your work and solve for the distance using additive inverses.

hotcheetos:

Noobsabotes:

the point 0 is positive 3

Noobsabotes:

no I got the answer wrong I think I did it let me redo it

Noobsabotes:

1.5+.5=2

Noobsabotes:

yea it 2

hotcheetos:

thank you so much

Noobsabotes:

np It don't take much to help a friend

surjithayer:

distance AB=(0.5-(-1.5))=0.5+1.5=2

jazzixlin:

yum